# Remembering Nino Boccara (1931–2018)

### 13 Dec 2023 v1

@@ 3 Statistical physics, 1960–1989
The papers on Ising model and its various generalizations and modifications form a substantial body of work, with over 20 publications in the period 1982-1989. They are mostly written in collaboration with A. Benyoussef (Faculté des sciences de Rabat, Morocco) and after 1986 also with several other coauthors. Some of these papers investigate very complicated models, for example, in the system considered in [1986PhRvB..34.7775B], the parameter space is six dimensional and the renormalization group tranformation which the authors use has 69 fixed points describing a large variety of critical behaviors.
@@ 4 Complex systems, 1989-2010
Acknowledgements I wish to thank Nazim Fatès and Anna Lawniczak for ecourangement to write this article, Franco Bagnoli for contributing his recollections and M. Catherine Kounelis for information about N.B. appointment at ESPCI. I am also very grateful to Nazim Fatès for reading and correcting the mansuscript.

### 04 Jan 2024 v2

@@ 3 Statistical physics, 1960–1989
The papers on Ising model and its various generalizations and modifications form a substantial body of work, with over 20 publications in the period 1982-1989. They are mostly written in collaboration with A. Benyoussef (Faculté des sciences de Rabat, Morocco) and after 1986 also with several other coauthors. Some of these papers investigate very complicated models, for example, in the system considered in [1986PhRvB..34.7775B], the parameter space is six dimensional and the renormalization group transformation which the authors use has 69 fixed points describing a large variety of critical behaviors.
@@ 4 Complex systems, 1989-2010
Acknowledgements I wish to thank Nazim Fatès and Anna Lawniczak for encouragement to write this article, Franco Bagnoli for contributing his recollections and M. Catherine Kounelis for information about N.B. appointment at ESPCI. I am also very grateful to Nazim Fatès for reading and correcting the manuscript.